OpenStudy (anonymous):

calculate the sum of 1/5 and 1/4

OpenStudy (anonymous):

to add fractions, they need to have the same denominator. so we will "force it" so. we want to find a common multiple of both denominators. taking 5x4 = 20 always works, but sometimes there are lower multiples. however not this case. we will make the denominators the same using a little math trick. notice that we technically multiplying by one. \[\frac{ 1 }{ 5 }+\frac{ 1 }{ 4 } = \frac{ 1 }{ 5 }\left( \frac{ 4 }{ 4 } \right) + \frac{ 1 }{ 4 }\left( \frac{ 5 }{ 5 } \right) = \frac{ 4 }{ 20 } + \frac{ 5 }{ 20 } = \frac{ 4+5 }{ 20 } = \frac{ 9 }{ 20 }\]

OpenStudy (anonymous):

so i multiplied both fractions by fractions which will make their denominators the same. and anything/itself = 1

OpenStudy (anonymous):

thanks I understand now

OpenStudy (anonymous):

yay :)